OpenStudy (anonymous):

x^2 2x _____ + _____ = ? x+y x+y

OpenStudy (kirbykirby):

you have the same denominator, so you can just combine the numerators together.. like if you had: \[ \frac{3}{5}+\frac{4}{5}=\frac{3+4}{5}=\frac{7}{5}\] Here, your "5" is like the "x+y" in your expression.. so you can just keep the denominator as "x+y" and just add the 2 numerators together.